# An evaluation of LLMs and Google Translate for translation of selected Indian languages via sentiment and semantic analyses

### v2

@@ 3.1.3 Hindi: Tamas
Tamas is a novel by Bhisham Sahni [92] written in Hindi that describes the implication of the division of India in the aftermath of the partition of 1947 [93]. More specifically, in a solemn and heart-wrenching manner, it describes the 1947 riots. The novel is unique for its unusual narrative that does not depict the conflict through the life of a certain character alone; instead, it spreads a wide vision of conflict from the insides of various people of different races, political affiliations and religious beliefs. The story covers the riots as a result, which marked the British Partition of India. The division aimed to separate the major Hindu and Muslim populations, into modern-day India and Pakistan, which ultimately resulted in bloody violence and displacement [94]. In 1988, Govind Nihalani adapted the book into a television series  [^4] Tamas: https://www.imdb.com/title/tt14751340/ starring Om Puri, which won multiple National Film Awards. The text features 21 chapters, and we will use a selected chapter for our study.
3.2 Data processing
@@
We obtained prominent translations of the respective texts in English and converted each of the selected texts from printable document format (PDF). We used the Natural Language Processing Toolkit (NLTK)  [^5] https://www.nltk.org/ to remove undesired characters that can appear due to conversion. Note that we obtain both the original and the translation of the respective texts and the original text is written in Devanagari script in the case of Sanskrit (Bhagavad Gita) and Hindi (Tamas) and the Telugu (Maha Prasthanam), respectively. The Telugu script features 54 letters - 16 vowels, 3 vowel modifiers, and 35 consonants. The Devanagari script is one of the official scripts of India and Nepal composed of 48 primary characters, including 14 vowels and 34 consonants [95].
@@
The process for the sentiment and semantic analysis of the chosen books, i.e. Bhagavad Gita, Maha Prasthanam and Tamas, involves several stages. Tokenisation refers to the process of breaking down a piece of text, such as a sentence or a paragraph, into individual words or “tokens.” Therefore, we begin with tokenisation, text cleaning and pre-processing which marks the initial phase of data extraction and processing. Then they were tokenised by verses for Bhagavad Gita and by poems for Maha Prasthanam. In the case of Tamas, we focused on sentiments classified based on small paragraphs. After the tokenised text was translated using different models. This approach allows a closer look at the text, indenting analysis of narrative components, the themes and the mood presented in each section.
3.3 Neural translation models
3.3.1 Google Translate
@@
Google Translate [96] marked a significant advancement in machine translation using Google Neural Machine Translation [97], leveraging statistical methods and later incorporating deep learning models [98]. Google Translate supports translation between over 100 languages, ranging from widely spoken languages like English, Spanish, and Chinese to less common ones such as Icelandic and Maori[99].Some studies have reviewed the effectiveness of Google Translate, such as translation in medical communication [100] and Sanskrit to English translation [74]. Tsai et al. [101] demonstrated that Google Translate English texts presented a number of components of significantly higher writing quality with fewer mistakes in spelling and grammar and fewer errors per word than its counterparts [101]. Prates et al. [102] studied gender biases and showed that Google Translate yields male defaults much more frequently than what would be indicated from demographic data alone. These studies motivate the comparison of Google  Translate with emerging LLMs, including GPT and Gemini.
3.3.2 GPT models
@@
ChatGPT  [^6] ChatGPT: https://openai.com/chatgpt is a highly sophisticated LLM [103] that revolutionised NLP with chatbot applications that include query answering systems, language translation, and tasks such as text summarisation, literature review and essay writing along with poems, prose, and songs [104]. GPT models use large amounts of data that include the Wikipedia corpus [105]. However, contrary to their ability to address various real-world tasks, the accuracy can be varied depending on how generic or discerning the input is leading to adequate or off-topic responses [106, 33]. ChatGPT has been constantly updating its features to support more languages and widely adopted by people from different regions of the world. Furthermore, GPT-4o is the advanced version of ChatGPT which is accessible through ChatGPT Plus which has a monthly subscription fee and has access to information up to December 2023. It also can search for real-time data. GPT-4o is multimodal, i.e. it can utilise data that includes text, speech and images. ChatGPT and GPT-4o leverage supervised learning and reinforcement learning from human feedback. ChatGPT supports more than 50 languages which are among them; English, French, German, Spanish, Chinese, Japanese and Arabic as well as other more doable languages [107]. We will evaluate both Chat-GPT and GPT-4 for translation of selected languages in our study.
@@
Gemini has been developed by Google [108] as a GPT-based LLM that [109] that uses English as the main language. The latest update (Gemini 3) incorporates more language versions: Italian, Portuguese and Turkish ensuring the service availability in over 70 countries. Gemini improves on the deficiencies of GPT-3 in terms of contextual understanding using a much larger text corpus and yielding more lengthy and well-structured narratives. However, the performance of Gemini is similar to ChatGPT and more research is needed to compare their performance for various tasks. In our study, we will specifically assess the potential of Gemini in the translation of Indian texts when compared to other GPT models and Google Translate.
@@
GPT-based models have been criticised for lack of accountability, transparency and expandability [110]. They can provide plausible correct answers, which at times turn out to be incorrect with expert validation. They also have a bias due to the training data type and source. Although these models have been recently introduced, studies about their strengths and weaknesses are being reported, eg. Lee et al. [111] reported that Gemini defeated GPT-4o in educational settings, where the findings suggest GPT-4o’s superior capability in handling complex multimodal educational tasks. Furthermore, Borji and Mohammadian [112] reported that GPT-4 emerged as the top-performing chatbot across almost all categories. These categories include reasoning, logic, facts, coding, bias, language, and humour. Gemini is less correlated with other models, while ChatGPT and GPT-4 are highly correlated in terms of their responses.
3.4 Semantic and Sentiment Analysis
@@
MPNet [113] is an advanced BERT model that is capable of semantic analysis while generating sentence embeddings. MPNet employs permuted language modelling, (unlike masked language modelling in BERT) and uses auxiliary position information as input and thus reduces the position discrepancy, enabling the model to utilise the full sentence. Hence, the sentences are assigned vector representations that capture the semantic meaning of the sentences [114]. These can be then compared to understand the similitude or dissimilarity between sentences. Therefore, MPNet has been specifically helpful for semantic search, which finds sentences that have a similar semantic meaning about a fixed query [115].
@@ 3.5 Framework
We first begin (Step 1) by processing the text in the electronic copies of the Bhagavad Gita (Sanskrit), Tamas (Hindi) and Maha Prasthanam (Telugu) obtained in their original script (Devanagari and Telugu). We split the texts into slokas (verses) in the case of Bhagavad Gita and sentences in the case of Tamas and Maha Prasthanam ordered by chapters after preprocessing. We also obtain expert translations of the texts, as shown in Table 2.
@@
Table 2: Details about authorship, date of publication and translation for the selected texts. Note that the original author of the Bhagavad Gita is unknown and its actual date of composition is difficult to access since it has been orally transmitted for centuries/millenniums.
@@
In Step 3, we provide translation vocabulary analysis using trigram analysis and compare the translations of the respective texts obtained by the LLMs and Google Translate. We tokenise the text for sentiment analysis using the BERT model [119]. We refine the BERT model using SenWave data for multi-label sentiment classification that can detect multiple emotions for a single text, which can be used for a deeper understanding of translation sentiments. The SenWave dataset has been used in our earlier studies [65, 74] for sentiment analysis of translations.
In Step 4, we use the MPNet model for semantic analysis and obtain a cosine similarity score between the expert translations and the translations. We then apply keyword extraction (Step 5) using KeyBERT which will extract the keywords from the text which can be analysed to know which keywords are dominant in various translations and we also extract the trigrams using KeyBERT.
@@
4 Results
4.1 Data Analysis
@@
As seen in Table 3, Google Translate strongly prefers theological terms, with high repetition of phrases (’supreme’, ’personality’, ’godhead’), suggesting a literal but somewhat redundant approach. Gemini focuses heavily on divine references, frequently using words such as ’krishna’ and ’lord,’ which portrays a devotional tone. GPT-3.5 and GPT-4o provide a more balanced interpretation, incorporating philosophical (’goodness’, ’passion’, ’ignorance’) and duty-related concepts (’perform’, ’prescribed’, ’duties’) that are associated with the philosophy of karma which is central to the Bhagavad Gita. In contrast, human expert annotations emphasise dialogue and context, using phrases such as (’lord’, ’shri’, ’krishna’) and (’arjuna’, ’asked’, ’lord’), which reflects the narrative style.
The trigram analysis for translations of Maha Prasthanam from Telugu-English (Table 4) reveals distinct patterns across different models and human expert translations. Google Translate heavily favours repetition, particularly with phrases (’forward’, ’forward’, ’forward’), indicating a mechanical translation approach. GPT-3.5 and GPT-4o provide more varied outputs, with a mix of movement-related words (’step’, ’forward’, ’step’) and (’come’, ’forward’, ’come’), though some trigrams appear redundant. Gemini emphasises directional movement, frequently using ’push’ and ’forward,’ which aligns with the theme of progress. In contrast, the human expert translation preserves poetic and philosophical depth, using phrases (’forward’, ’march’, ’onward’) and (’manifold’, ’divine’, ’lights’), which better capture the literary essence. This comparison highlights that while the different models can extract structure and meaning, the human expert translation retains the deeper poetic and contextual nuances of the original text.
In the case of Tamas translation from Hindi-English (Table 5) GPT-3.5 and GPT-4o generate a mix of descriptive and action-based trigrams, such as (”pigs”, ”roam”, ”around”) and (”walking”, ”along”, ”wall”), capturing movement and scene details. Google Translate produces a mix of mechanically translated phrases, often repeating phrases (”opened”, ”door”, ”closet”), showing a more direct and structured translation. We observe the common trigram, (”flickering”, ”light”, ”lamp”) in GPT-3.5, GPT-4o, and Gemini, indicating a shared interpretation of the text’s imagery. However, in the Google Translate output, the phrases (”lamp”, ”started”, and ”blinking”) appeared instead, reflecting a more literal translation from the original Hindi text.
@@
Table 3: Trigrams Analysis for Bhagavad Gita translation from Sanskrit to English with LLMs, Google Translate and Human Expert.
@@
Table 4: Trigrams Analysis for Maha Prasthanam translation from Telugu to English with LLMs, Google Translate and Human Expert.
@@
Table 5: Trigrams Analysis for Tamas translation from Hindi to English with LLMs, Google Translate and Human Expert.
4.2 Sentiment Analysis
@@
Table 6: Sentiments Counts(Sanskrit)
@@
Table 7: Sentiments Counts(Telugu)
@@
Table 8: Sentiments Counts(Hindi)
We further present results from sentiment analysis of translations by LLMs (Gemini and GPT models) for Sanskrit, Hindi and Telugu and compare the sentiments with human-expert translation. In this way, we can compare if the different translations express similar sentiments along with the varied priority of the sentiments. In the case of the Bhagavad Gita translations from Sanskrit-English (Figure 2), we find that the respective models are visibly different in capturing various sentiments. We note that the ”joking” sentiment refers to humour and this was detected by human expert translators such as Mahatma Gandhi, Sri Purohit Swami and Eknath Easwaren in a study by Chandra and Kulkarni [65]. Google Translate produced the highest number of ”Optimistic” (293) and ”Empathetic” (7) sentiments for Sanskrit-English translation (Figure 2), suggesting that its translation is more positively expressed sentiments than others. However, Google Translate also generated the highest number of ”Pessimistic” (274) sentiments, indicating inconsistency in emotional tone. GPT-4o and the Human Expert had similar distributions across most sentiments, suggesting that GPT-4o’s translation is more aligned with human understanding. GPT-3.5 and GPT-4o had comparable sentiment counts, with GPT-4o slightly increasing the presence of emotions such as ”Pessimistic” and ”Denial.”
@@
Figure 3 presents the sentiments detected for the Telugu-English translations; GPT-4o and Gemini had the highest ”Pessimistic” counts (35 and 34), while the Human Expert had a similar count (33). ”Optimistic” sentiment was highest in the Human Expert’s translation (19), followed by GPT-3.5 (17). ”Anxious” and ”Sad” counts were similar across models, with Gemini and GPT-3.5 showing the highest anxiety (10 each). Google Translate had the most ”Empathetic” (3) and ”Denial” (2) sentiments, while all other models had zero in these categories. Google Translate also had the highest ”Annoyed” sentiment (14), whereas Gemini had the lowest (4). GPT-4o and GPT-3.5 showed nearly identical distributions in the ”Annoyed” and ”Sad” categories. Overall, GPT-4o and Gemini leaned towards pessimism, while Google Translate introduced more emotional variation. The Human Expert’s sentiment distribution remained more balanced, closely aligning with GPT models in most categories.
@@
Finally, in the case of the Hindi-English translations (Figure 4, Gemini produced the highest number of ”Optimistic” (9) sentiments, suggesting that its translations may lean toward a slightly more positive tone. In contrast, Google Translate had the lowest count (4), indicating a more neutral or less expressive output. Interestingly, the Human Expert showed the highest number of ”Pessimistic” (26) sentiments, implying that human translations might retain more nuance and emotional depth from the original text. Google Translate also registered the highest number of ”Sad” (22) and ”Annoyed” (30) sentiments, indicating a potential shift in tone that makes the text sound more negative. While GPT-4o and the Human Expert had similar ”Annoyed” counts (26 and 27), Gemini remained slightly lower at 23. ”Thankful” sentiments were nearly absent across all translations, appearing only once in Gemini’s output. Similarly, ”Empathetic” emotions were rare, with the Human Expert producing the highest count (2), indicating that models struggle to preserve subtle expressions of empathy. ”Denial” was also relatively low across all translations, with GPT-4o having the highest count (7), possibly due to variations in phrasing and interpretation.
@@
Table 9: Mean and Standard Deviation of TextBlob Sentiment Polarity Score for Different Models.
@@
Figure 6 presents the Telugu-English translations, overall GPT-4o demonstrates the highest alignment with human expert translations, maintaining relatively consistent polarity scores across sentence counts but deviating towards the end (plus 50 sentences). Gemini follows closely, though it exhibits sharper fluctuations, particularly in the mid-range sentences. GPT-3.5 and Google Translate, however, show significant deviations from the human expert, with Google Translate exhibiting the most erratic behaviour, especially towards the end of the sentence range (plus 50 sentences). Taking a lot at the sentences towards the end, where we observe a higher deviation since more figurative elements have been used in the sentences which might have led to inaccurate translation, and translation that lacked the sentiment of the actual sentence. For example (Sentence 47): “On the blade of the dagger, it is like a drop of perspiration” this was the translation by the human expert, and on the other hand, “It looks like a drop of blood on the sword handle”, it is seen that model has translated the sentence when it suggested the tension, fragility by the image of the drop of perspiration.
@@
4.3 Semantic Analysis
We use the MPNet-Base model[46] providing semantic analysis for comparing different translations by LLMs to a human expert. We evaluate the translations by GPT-4o, GPT-3.5, Google Translate and Gemini using the cosine similarity scores (from 0 to 1), where 1 implies a perfect semantic match. We first provide semantic similarity analysis of the respective language translation models across Sanskrit to English translations as shown in Table 11. We observe that GPT-4o achieves the highest average similarity of 0.686, followed by GPT-3.5 at 0.665 and Gemini at 0.660. Although GPT-3.5 attains the highest peak similarity of 0.979, GPT-4o maintains a more consistent performance. Additionally, Google Translate lags with an average similarity of 0.603. Table 10 presents further details with chapter-wise similarity scores where GPT-4o provides the best scores.
In the case of Telugu-English translation (Table 12, Gemini leads with an average similarity of 0.726, outperforming all other models. GPT-4o follows with 0.704, while GPT-3.5 records a lower score of 0.583, showing lower similarity to the human expert translations. Google Translate performs relatively better in Telugu compared to Sanskrit, securing an average similarity of 0.656 but still trailing behind the leading models.
Finally, in the case of Hindi-English translations (Table 11), GPT-4o showed the highest similarity to human translations, with an average cosine similarity of 0.754 and a peak of 0.927, indicating strong alignment with human understanding. Gemini followed with 0.739, performing slightly better than GPT-3.5 (0.711) but lower than GPT-4o. Google Translate had the lowest average similarity (0.684), suggesting that its translations were less accurate overall. However, its highest similarity score (0.916) was close to that of other models, meaning it could sometimes produce accurate results. The lowest similarity scores highlight consistency differences—GPT-4o had the highest minimum similarity (0.406), while GPT-3.5 dropped to 0.317, showing that its translations were sometimes less reliable.

### v3

@@ III-A3 Hindi: Tamas
Tamas is a novel by Bhisham Sahni [92] written in Hindi that describes the implications of the division of India in the aftermath of the partition of 1947 [93]. More specifically, in a solemn and heart-wrenching manner, it describes the 1947 riots. The novel is unique for its unusual narrative that does not depict the conflict through the life of a certain character alone; instead, it spreads a wide vision of conflict from the insides of various people of different races, political affiliations and religious beliefs. The story covers the riots as a result, which marked the British Partition of India. The division aimed to separate the major Hindu and Muslim populations into modern-day India and Pakistan, which ultimately resulted in bloody violence and displacement [94]. In 1988, Govind Nihalani adapted the book into a television series  [^4] Tamas: https://www.imdb.com/title/tt14751340/ starring Om Puri, which won multiple National Film Awards. The text features 21 chapters, and we will use a selected chapter for our study.
III-B Data processing
@@
We obtained prominent translations of the respective texts in English and converted each of the selected texts in a Printable Document Format (PDF). In the case of the Sanskrit language (Bhagavad Gita), we used the entire text, and in the case of Telugu and Hindi, we used selected chapters that were openly available for usage. We used the Natural Language Processing Toolkit (NLTK)  [^5] https://www.nltk.org/ to remove undesired characters that can appear due to conversion. Note that we obtain both the original and the translation of the respective texts, and the original text is written in Devanagari script in the case of Sanskrit (Bhagavad Gita) and Hindi (Tamas) and the Telugu (Maha Prasthanam), respectively. The Telugu script features 54 letters - 16 vowels, 3 vowel modifiers, and 35 consonants. The Devanagari script is one of the official scripts of India and Nepal, composed of 48 primary characters, including 14 vowels and 34 consonants [95].
@@
The process for the sentiment and semantic analysis of the chosen books, i.e. Bhagavad Gita, Maha Prasthanam and Tamas, involves several stages. Tokenisation refers to the process of breaking down a piece of text, such as a sentence or a paragraph, into individual words or “tokens.” Therefore, we begin with tokenisation, text cleaning and pre-processing, which marks the initial phase of data extraction and processing. Then they were tokenised by verses for the Bhagavad Gita and by poems for the Maha Prasthanam. In the case of Tamas, we focused on sentiments classified based on small paragraphs. Afterwards, the tokenised text was translated using different machine translation models (LLMs).
III-C Neural translation models
III-C1 Google Translate
@@
Google Translate [96] marked a significant advancement in machine translation using Google Neural Machine Translation (NMT) [97], leveraging statistical methods and later incorporating deep learning models [98]. Google Translate supports translation between over 100 languages, ranging from widely spoken languages such as English, Spanish, and Chinese to less common ones such as Icelandic and Maori[99]. Some studies have reviewed the effectiveness of Google Translate, including translations in medical communication [100] and Sanskrit to English translation [74]. Tsai et al. [101] demonstrated that Google Translate English texts presented several components of significantly higher writing quality with fewer mistakes in spelling and grammar and fewer errors per word than its counterparts [101]. Prates et al. [102] studied gender biases and showed that Google Translate yields male defaults much more frequently than what would be indicated from demographic data alone. These studies motivate the comparison of Google  Translate with emerging LLMs, including GPT and Gemini.
III-C2 GPT models
@@
ChatGPT  [^6] ChatGPT: https://openai.com/chatgpt is a highly sophisticated LLM [103] that revolutionised NLP with chatbot applications that include query answering systems, language translation, and tasks such as text summarisation, literature review and essay writing, along with poems, prose, and songs [104]. GPT models use large amounts of data that include the Wikipedia corpus [105]. However, contrary to their ability to address various real-world tasks, the accuracy can be varied depending on how generic or discerning the input is leading to adequate or off-topic responses [106, 33]. ChatGPT has been constantly updating its features to support more languages and has been widely adopted by people from different regions of the world. Furthermore, GPT-4o is the advanced version of ChatGPT which is accessible through ChatGPT Plus, which has a monthly subscription fee and has access to information up to December 2023, and can also search for real-time data. GPT-4o is multimodal, i.e. it can utilise data that includes text, speech and images. ChatGPT and GPT-4o leverage supervised learning and reinforcement learning from human feedback. ChatGPT supports more than 50 languages which are including English, French, German, Spanish, Chinese, Japanese and Arabic, as well as other more doable languages [107]. We will evaluate both Chat-GPT (GPT-3.5) and GPT-4o for translation of selected languages in our study.
@@
Gemini has been developed by Google [108] as a GPT-based LLM that [109] uses English as the main language. The latest update (Gemini 3) incorporates more language versions: Italian, Portuguese and Turkish, ensuring the service availability in over 70 countries. Gemini improved on the deficiencies of GPT-3 in terms of contextual understanding using a much larger text corpus and yielding more lengthy and well-structured narratives. However, the performance of Gemini is similar to ChatGPT, and more research is needed to compare their performance for various tasks. In our study, we will specifically assess the potential of Gemini in the translation of Indian texts, when compared to other LLMs and Google Translate.
@@
GPT-based models have been criticised for a lack of accountability, transparency and expandability [110]. They can provide plausible, correct answers, which at times turn out to be incorrect with expert validation. They also have a bias due to the training data type and source. Although these models have been recently introduced, studies about their strengths and weaknesses are being reported, eg. Lee et al. [111] reported that Gemini defeated GPT-4o in educational settings, where the findings suggest GPT-4o’s superior capability in handling complex multimodal educational tasks. Furthermore, Borji and Mohammadian [112] reported that GPT-4 emerged as the top-performing chatbot across almost all categories. These categories include reasoning, logic, facts, coding, bias, language, and humour. Gemini is less correlated with other models, while ChatGPT and GPT-4 are highly correlated in terms of their responses.
III-D Semantic and Sentiment Analysis
@@
MPNet [113] is an advanced BERT model that is capable of semantic analysis while generating sentence embeddings. MPNet employs permuted language modelling, (unlike Masked Language Modelling in BERT) and uses auxiliary position information as input and thus reducing the position discrepancy, enabling the model to utilise the full sentence. Hence, the sentences are assigned vector representations that capture the semantic meaning of the sentences [114]. These can be then compared to understand the similitude or dissimilarity between sentences. Therefore, MPNet has been specifically helpful for semantic search, which finds sentences that have a similar semantic meaning about a fixed query [115].
@@ III-E Framework
We first begin (Step 1) by processing the text in the electronic copies of the Bhagavad Gita (Sanskrit), Tamas (Hindi) and Maha Prasthanam (Telugu) obtained in their original script (Devanagari and Telugu). We split the texts into slokas (verses) in the case of Bhagavad Gita and sentences in the case of Tamas and Maha Prasthanam ordered by chapters after preprocessing. We also obtain expert translations of the texts, as shown in Table II.
@@
TABLE II: Details about authorship, date of publication and translation for the selected texts. Note that the original author of the Bhagavad Gita is unknown, and its actual date of composition is difficult to access since it has been orally transmitted for centuries/millennia.
@@
In Step 3, we provide translation vocabulary analysis using trigram analysis and compare the translations of the respective texts obtained by the LLMs and Google Translate. We tokenise the text for sentiment analysis using the BERT model [119]. We refine the BERT model using SenWave data [120] for multi-label sentiment classification that can detect multiple emotions for a single text, which can be used for a deeper understanding of translation sentiments. The main motivation of using BERT-based SenWave model for sentiment analysis is due to its multi-label classification nature, where more than one sentiment can be detected in a sentence. Moreover, it has been used effectively in previous research [65, 74] for sentiment analysis in the scope of machine translation. However, we note the limitation that the SenWave dataset is X (Twitter) dataset, while the applications in this study consists of philosophical and sacred texts and novellas.
In Step 4, we use the MPNet model for semantic analysis and obtain a cosine similarity score between the expert translations and the translations. We then apply keyword extraction (Step 5) using KeyBERT, which will extract the keywords from the text, which can be analysed to know which keywords are dominant in various translations, and we also extract the trigrams using KeyBERT.
@@
IV Results
IV-A Data Analysis
@@
As seen in Table III, Google Translate strongly prefers a translation using theological terms, with high repetition of phrases (’supreme’, ’personality’, ’godhead’), suggesting a literal but somewhat redundant approach. Gemini focuses heavily on divine references, frequently using words such as ’krishna’ and ’lord,’ which portrays a devotional tone. GPT-3.5 and GPT-4o provide a more balanced interpretation, incorporating philosophical (’goodness’, ’passion’, ’ignorance’) and duty-related concepts (’perform’, ’prescribed’, ’duties’) that are associated with the philosophy of karma, which is central to the Bhagavad Gita. In contrast, human expert annotations emphasise dialogue and context, using phrases such as (’lord’, ’shri’, ’krishna’) and (’arjuna’, ’asked’, ’lord’), which reflect the narrative style.
The trigram analysis for translations of Maha Prasthanam from Telugu-English (Table IV) reveals distinct patterns across different models and human expert translations. Google Translate heavily favours repetition, particularly with phrases (’forward’, ’forward’, ’forward’), indicating a mechanical translation approach. GPT-3.5 and GPT-4o provide more varied outputs, with a mix of movement-related words (’step’, ’forward’, ’step’) and (’come’, ’forward’, ’come’), though some trigrams appear redundant. Gemini emphasises directional movement, frequently using ’push’ and ’forward,’ which aligns with the theme of progress. In contrast, the human expert translation preserves poetic and philosophical depth, using phrases (’forward’, ’march’, ’onward’) and (’manifold’, ’divine’, ’lights’), which better capture the literary essence. This comparison highlights that while the different models can extract structure and meaning, the human expert translation retains the deeper poetic and contextual nuances of the original text.
In the case of Tamas translation from Hindi-English (Table V) GPT-3.5 and GPT-4o generate a mix of descriptive and action-based trigrams, such as (”pigs”, ”roam”, ”around”) and (”walking”, ”along”, ”wall”), capturing movement and scene details. Google Translate produces a mix of mechanically translated phrases, often repeating phrases (”opened”, ”door”, ”closet”), showing a more direct and structured translation. We observe the common trigram (”flickering”, ”light”, ”lamp”) in GPT-3.5, GPT-4o, and Gemini, indicating a shared interpretation of the text’s imagery. However, in the Google Translate output, the phrases (”lamp”, ”started”, and ”blinking”) appeared instead, reflecting a more literal translation from the original Hindi text.
@@
TABLE III: Trigrams Analysis for Bhagavad Gita translation from Sanskrit to English with LLMs, Google Translate and Human Expert.
@@
TABLE IV: Trigrams Analysis for Maha Prasthanam translation from Telugu to English with LLMs, Google Translate and Human Expert.
@@
TABLE V: Trigrams Analysis for Tamas translation from Hindi to English with LLMs, Google Translate and Human Expert.
IV-B Sentiment Analysis
@@
TABLE VI: Sentiments Counts(Sanskrit)
@@
TABLE VII: Sentiments Counts(Telugu)
@@
TABLE VIII: Sentiments Counts(Hindi)
We further present results from sentiment analysis of translations by LLMs (Gemini and GPT models) for Sanskrit, Hindi and Telugu and compare the sentiments with human-expert translation. In this way, we can compare whether the different translations express similar sentiments along with the varied priority of the sentiments. In the case of the Bhagavad Gita translations from Sanskrit-English (Figure 2), we find that the respective models are visibly different in capturing various sentiments. We note that the ”joking” sentiment refers to humour and this was detected by human expert translators such as Mahatma Gandhi, Sri Purohit Swami and Eknath Easwaren in a study by Chandra and Kulkarni [65]. Google Translate produced the highest number of ”Optimistic” (293) and ”Empathetic” (7) sentiments for Sanskrit-English translation (Figure 2), suggesting that its translation is more positively expressed sentiments than others. However, Google Translate also generated the highest number of ”Pessimistic” (274) sentiments, indicating inconsistency in emotional tone. GPT-4o and the Human Expert had similar distributions across most sentiments, suggesting that GPT-4o’s translation is more aligned with human understanding. GPT-3.5 and GPT-4o had comparable sentiment counts, with GPT-4o slightly increasing the presence of emotions such as ”Pessimistic” and ”Denial.”
@@
Figure 3 presents the sentiments detected for the Telugu-English translations; GPT-4o and Gemini had the highest ’Pessimistic’ sentiment counts (35 and 34), while the Human Expert has a similar count (33). We notice that the ’Optimistic’ sentiment is highest in the Human Expert’s translation (19), followed by GPT-3.5 (17). Furthermore, the ’Anxious’ and ’Sad’ counts are similar across models, with Gemini and GPT-3.5 showing the highest anxiety (10 each). Google Translate has the most ’Empathetic’ (3) and ’Denial’ (2) sentiments, while all other models had zero in these categories. Google Translate also had the highest ’Annoyed’ sentiment (14), whereas Gemini had the lowest (4). GPT-4o and GPT-3.5 showed nearly identical distributions in the ”Annoyed” and ”Sad” categories. Overall, GPT-4o and Gemini leaned towards pessimism, while Google Translate introduced more emotional variation. The Human Expert’s sentiment distribution remains more balanced, closely aligning with LLMs in most categories.
@@
Finally, in the case of the Hindi-English translations (Figure 4, Gemini produced the highest number of ”Optimistic” (9) sentiments, suggesting that its translations may lean toward a slightly more positive tone. In contrast, Google Translate has the lowest count (4), indicating a more neutral or less expressive output. Interestingly, the Human Expert showed the highest number of ”Pessimistic” (26) sentiments, implying that human translations might retain more nuance and emotional depth from the original text. Google Translate also featured the highest number of ”Sad” (22) and ’Annoyed’ (30) sentiments, indicating a potential shift in tone that makes the text sound more negative. Although GPT-4o and the Human Expert had similar ’Annoyed’ counts (26 and 27), Gemini remained slightly lower at 23. We find that ’Thankful’ is nearly absent from all translations, appearing only once in Gemini’s output. Similarly, we find the ’Empathetic’ emotion rarely expressed, indicating that LLMs struggle to preserve subtle expressions of empathy. In comparison, we find that the Human Expert producing an ’Empathetic’ emotion count of 2. ’Denial’ is also relatively low across all translations, with GPT-4o having the highest count (7), possibly due to variations in phrasing and interpretation.
@@
TABLE IX: Mean and Standard Deviation of TextBlob Sentiment Polarity Score for Different Models.
@@
Figure 6 presents the Telugu-English translations; overall GPT-4o demonstrates the highest alignment with human expert translations, maintaining relatively consistent polarity scores across sentence counts but deviating towards the end (plus 50 sentences). Gemini follows closely, though it exhibits sharper fluctuations, particularly in the mid-range sentences. GPT-3.5 and Google Translate show significant deviations from the human expert, with Google Translate exhibiting the most erratic behaviour, especially towards the end of the sentence range (plus 50 sentences). We review the sentences towards the end, where we observe a higher deviation since more figurative elements have been used in the sentences, which might have led to inaccurate translation, and a translation that lacked the sentiment of the actual sentence. For example, Sentence 47: “On the blade of the dagger, it is like a drop of perspiration” this was the translation by the human expert, and on the other hand, “It looks like a drop of blood on the sword handle”. We find that the model has translated the sentence when it suggested the tension, fragility by the image of the drop of perspiration.
@@
IV-C Semantic Analysis
We use the MPNet-Base model[46], providing semantic analysis for comparing different translations by LLMs to a human expert. We evaluate the translations by GPT-4o, GPT-3.5, Google Translate and Gemini using the cosine similarity scores (from 0 to 1), where 1 implies a perfect semantic match. We first provide semantic similarity analysis of the respective language translation models across Sanskrit to English translations as shown in Table XI. We observe that GPT-4o achieves the highest average similarity of 0.686, followed by GPT-3.5 at 0.665 and Gemini at 0.660. Although GPT-3.5 attains the highest peak similarity of 0.979, GPT-4o maintains a more consistent performance. Additionally, Google Translate lags with an average similarity of 0.603. Table X presents further details with chapter-wise similarity scores, where GPT-4o provides the best scores. Table XI presents further details where the Sanskrit-English translation has been compared.
In the case of Telugu-English translation (Table XII, Gemini leads with an average similarity of 0.726, outperforming all other models. GPT-4o follows with 0.704, while GPT-3.5 records a lower score of 0.583, showing lower similarity to the human expert translations. Google Translate performs relatively better in Telugu compared to Sanskrit, securing an average similarity of 0.656, but still trails behind the leading models.
Finally, in the case of Hindi-English translations (Table XIII) GPT-4o achieves the highest similarity to human translations, with an average cosine similarity of 0.754 and a peak of 0.927, indicating strong alignment with human understanding. Gemini followed with 0.739, performing slightly better than GPT-3.5 (0.711), but lower than GPT-4o. Google Translate has the lowest average similarity (0.684), suggesting that its translations are less accurate overall. However, its highest similarity score (0.916) was close to that of other models, meaning it could sometimes produce accurate results. The lowest similarity scores highlight consistency differences. GPT-4o has the highest minimum similarity (0.406), while GPT-3.5 dropped to 0.317, showing that its translations are sometimes less reliable.